# Quantification of the Self-Excited Emotion Dynamics in Online Interactions

### v1

@@ Methods
Note that Equation (5) represents the log-likelihood for a single video, $\log L^{e}=\log L^{e}_{k}$ where $k$ enumerates the videos in our data sample. We thus aggregate $L^{e}_{k}$ across all videos and estimate the model parameters by maximizing $\sum_{k}\log L^{e}_{k}$. The total number of parameters is $142|\mathcal{E}|+2=$ 14 per emotion. This amounts to a total of $|\mathcal{E}|\left(2|\mathcal{E}|+2\right)=84$ parameters to be estimated for $|\mathcal{E}|=6$ emotions. However, the parameters for each emotion can be estimated independently because there is no interdependence between the parameters across different emotion types. This gives rise to six independent fits of $14$ parameters each. To verify that the obtained fits are valid, we first test our model with synthetically generated data and obtain reliable performance (see SI Appendix). We also refer to the SI Appendix for additional robustness checks with respect to different parametrizations and subsets of emotions. To mitigate potential biases from Simpson’s paradox, where combining individual groups may lead to spurious trends, we estimate our model parameters using bootstrapped samples. We bootstrap our data $10$ times, each time sampling only 60% of all available videos. This allows us to estimate the parameters as averages across bootstrapped samples and we use standard deviations as error bars. The fitted parameters are shown in Figure 2.
@@ Appendix A Summary Statistics of YouTube Live Chat Data
We show some additional summary statistics of our dataset. The final sample of our data contains 92,412 live chat messages across 397 videos, adding up to 780 hours of content and covering 27 topics. The distribution of topics is shown in Figure A.1. The distribution of emotional events is shown in Figure A.2. We also show the distribution of the median time between each event arrival in Figure A.3. Since we label emotions non-exclusively, we show the co-occurrence patterns across 6 basic emotions in Figure A.4. In general, negative emotions have a higher tendency to co-occur with one another. In particular, anger and disgust have a relatively high co-occurrence.
@@ Appendix B Robustness Analysis on Video Influence
To examine the impact of video content on live chat discussions, we have parameterized and interpolated the prevalence of emotions in the video via sums of log-normal shapes (see Methods). Here, we test the robustness of our results with respect to changes in the log-normal parameters. We have defined the log-normal function with 2 parameters, the maximum, and the median. We assumed that the intensity of each transcript peaks $2$ seconds after appearance such that the maximum of the log-normal function lies at $2$ seconds. We further assumed that 50% of the emotion intensity for each transcript is manifested within $10$ seconds of transcript appearance such that the median is equal to $10$ seconds. Here we repeat our analysis while changing the log-normal function maximum and median by factors of 0.5 and 2 to test the robustness of our results. This effectively changes assumptions about the audience’s reaction time and memory of prior content. We observe in Figures 1(a), 1(b), 2(a), and 2(b) that the results with the different log-normal parameters are highly consistent with our results in the main text.
@@ Appendix C Dynamics of 4 Emotions
Our main analysis models the commonly studied 6 basic emotions [34]. Here we test the robustness of our results by modeling $|\mathcal{E}|=4$ emotions: joy, anger, disgust, and sadness, removing fear and surprise because they have small excitatory patterns and low overall intensity. We obtain a sample of 288,581 live chat messages across 1,405 videos. Reducing the dimensionality of the multivariate Hawkes process reduces the number of parameters to be estimated $|\mathcal{E}|\left(2|\mathcal{E}|+2\right)$ from 84 to 40.
We obtain our results for 4 emotions in Figures C.1 and C.2. We find that the dynamics of 4 emotions are qualitatively similar to our main results on 6 emotions (Figures 2 and 3 in the main text).
@@ Appendix E Dynamics of Live Chats with No Emotional Content
Table 2: The mean and standard deviation of estimated parameters for a univariate Hawkes Process of live chat messages with no emotion labels across 10 bootstrapped samples.
@@
Our main study shows that online emotions are characterized by excitatory patterns. Despite the constant feed of video content, the emotion dynamics evolve mostly spontaneously among users. To complement our analysis, we study the behavior of live chats with no emotional content. We train a transformer to assign emotion labels for live chats. We deem the live chat messages that were not assigned any emotion labels as containing no emotional content. This gives us 66,235 live chat messages with no emotional content across 232 videos. We use a univariate Hawkes process to describe the intensity of live chat messages with no emotion labels. The intensity $\lambda(t)$ at time $t$ during a given video is specified as
@@
The value of $S(t)$ is given by $S(t)=\sum_{\tau_{j}<t}s_{\tau_{j}}(t)$ where $\left\{\tau_{j}\right\}$ enumerates all times at which a subtitle first appears in the video. The values of $s_{\tau_{j}}(t)$ are computed according to Equation (3) from the main text with $\mu=2.3$ and $\sigma=1.3$. The term $s_{\tau_{j}}(t)$ represents the time-varying intensity arising from a video subscript appearing in the video at time $\tau_{j}$. Similarity, we assume that $\phi$ is exponentially decaying where $\phi(t)=\alpha\frac{1}{\gamma}e^{-\frac{1}{\gamma}t}$ with decay time $\gamma$. Table 2 shows the estimated parameters of Equation (1). We observe that live chat messages with no emotional content are nonetheless highly self-exciting, but with a shorter direct memory than emotional live chats. Different from our main result (Figure 3 in the main text), live chat messages with no emotional content are driven more equally by the video and user spontaneity as shown in Figure E.1. This shows that while live discussions largely evolve endogenously, live chats with no emotional content are much more affected by the videos. In contrast, emotions on online platforms are largely attributed to user expressions and interactions.

### v2

@@ Methods
Note that Equation (5) represents the log-likelihood for a single video, $\log L^{e}=\log L^{e}_{k}$ where $k$ enumerates the videos in our data sample. Thus, we aggregate $\log L^{e}_{k}$ across all videos and estimate the model parameters by maximizing $\sum_{k}\log L^{e}_{k}$. The total number of parameters is $142|\mathcal{E}|+2=$ 14 per emotion. This amounts to a total of $|\mathcal{E}|\left(2|\mathcal{E}|+2\right)=84$ parameters to estimate for $|\mathcal{E}|=6$ emotions. However, the parameters for each emotion can be estimated independently because there is no interdependence between the parameters between different types of emotions. This gives rise to six independent fits of $14$ parameters each. To verify that the obtained fits are valid, we first test our model with synthetically generated data and obtain reliable performance (see SI Appendix). We also refer to the SI Appendix for additional robustness checks with respect to different parameterizations and subsets of emotions. To mitigate potential biases, we estimate our model parameters using bootstrapped samples. We bootstrap our data $10$ times, each time sampling only 60% of all available videos. This allows us to estimate the parameters as averages across bootstrapped samples, and we use standard deviations as error bars. The fitted parameters are shown in Figure 2. These parameters are interpreted in the Results Section and discussed more qualitatively in the Discussion Section.
Results
Figure 2: Visualizations of estimated parameters $\nu^{e,f}$, $\mu^{e}_{0}$, $\alpha^{e,f}$, and $\gamma^{e}$ of the intensity defined by expression (1) from the maximization of the log-likelihood function (5). The parameters are fitted ten times, each time sampling 60% (238) of the total 397 videos at random. Values and error bars are then obtained as sample mean and standard deviations across all 10 fits, respectively. Entry $\alpha^{e,f}$ of the $\alpha$ matrix represents excitation from emotion $f$ to emotion $e$ in the live chat, while $\gamma^{e}$ represents the characteristic time-scale over which past emotions trigger new emotions $e$. The $\nu$ matrix demonstrates how emotions in the video trigger emotions in the live chat, while $\mu_{0}$ illustrates the spontaneous baseline intensity of spontaneous emotion expression.
Quantifying Exogenous and Endogenous Influences
Figure 3: (a) For each video, we calculate the average ratio of endogenous (exogenous) intensity to the total intensity across time. Plot (a) shows the distribution of these ratios across videos. We notice that emotions are predominantly triggered endogenously. (b) Same as plot (a) but for the ratio of spontaneous (video-influenced) intensity to total exogenous intensity.
Videos provide a shared experience among viewers, resulting in increased emotional intensities during live broadcast events [12]. However, it is unclear to what extent the emotional experience and amplification stem from the live videos themselves or the interactions among peers [12]. The Hawkes framework (1) allows us to answer this question by comparing exogenously and endogenously triggered events. The exogenous component, $\mu^{e}(t)=\mu^{e}_{0}+\mu_{1}^{e}(t)$, represents influences from video content and elsewhere. The endogenous component, $\sum_{f\in\mathcal{E}}\sum_{t_{j}^{f}<t}\phi^{e,f}\left(t-t_{j}^{f}\right)$, captures the influence of previous live chat messages and is therefore derived from interactions with peers. At any given time, we can determine the fraction of exo-influence as $R^{e}_{\text{exo}}(t)=\left.\mu^{e}(t)\right/\lambda^{e}(t)$ and the ratio of endo influence as $R^{e}_{\text{endo}}(t)=1-R^{e}_{\text{exo}}(t)$ where all quantities are calculated with the estimated parameters (shown in Figure 2 and further discussed below). Subsequently, we calculate the average exo- and endo-ratios, $\langle R^{e}_{\text{exo}}\rangle$ and $\langle R^{e}_{\text{endo}}\rangle$, of a given video by averaging over time. To put it differently, for each video, $\langle R^{e}_{\text{exo (exo)}}\rangle$ represents the mean portion of intensity for emotion $e$ that is elicited exogenously (endogenously). Figure 3 (left) shows the distribution of these quantities over all videos in our sample. We observe that despite the constant exogenous influence from video feeds, endogenous influence dominates the emotion dynamics across emotion types. This shows that emotional expressions in live discussions are disproportionately driven by social interactions with active participants rather than passive consumption of video content. In particular, we see that the emotion of joy has the highest ratio of endogenous influence, being 80% driven by prior participants. In other words, joy is four times more likely to be triggered endogenously than exogenously. In contrast, and according to common sense, the emotion of surprise is the most responsive to exogenous stimuli, where roughly 50% of the total intensity is triggered exogenously.
Similarly, we can calculate the ratio of purely exogenous spontaneous emotions relative to video-induced emotions, that is $R^{e}_{0}=\mu^{e}_{0}/\left(\mu^{e}_{0}+\mu^{e}_{1}(t)\right)$ and its complement $R^{e}_{1}=1-R^{e}_{0}.$ Figure 3 (right) shows the distribution of $\langle R^{e}_{0}\rangle$ versus $\langle R^{e}_{1}\rangle$ over the set of all videos. The rate of chats induced by videos is typically one-fourth the rate of spontaneous chats. We note, however, that there is significant heterogeneity across video categories (SI Appendix). In summary, this highlights that live chat discussions elicit emotional expressions largely due to social dynamics rather than simply watching video content. In the SI Appendix, we further support these findings with additional robustness checks, specifically focusing on outcomes when fitting only selected subsets of emotions or video categories.
Emotion Contagion and Interaction Dynamics
We now interpret the estimated parameters of Equation (1) visualized in Figure 2. Here, we focus more on the quantitative aspects of our fit, whereas a qualitative interpretation is given in the following Discussion section. Recall that $\alpha^{e,f}$ (Figure 2, left) represents the rate at which chat-induced emotion $f$ leads to emotion $e$ within the chat (endo). Conversely, $\nu^{e,f}$ (Figure 2, right) signifies the rate at which video-induced emotion $f$ induces emotion $e$ within the chat (exo). In the preceding section, we observed that most live chat messages originate endogenously, which means that, on average, $\alpha^{e,f}$ significantly exceeds $\nu^{e,f}$. Consequently, we shall concentrate our analysis on $\alpha^{e,f}$.
A key quantity is the spectral radius of the branching ratio matrix $\alpha^{e,f}$, denoted $\rho$, which provides a global measure of self-excitation and cross-excitation. Mathematically, the spectral radius is defined as the dominant eigenvalue of the matrix $\alpha^{e,f}$, reflecting the overall rate at which events trigger new events. It captures the extent to which activity is self-sustaining or dependent on external input. If $\rho>1$, the process is supercritical, which means that each event generates, on average, more than one new event, leading to an exponential explosion of activity. This scenario is generally not realistic, as unbounded growth is unsustainable in most real-world systems, except in rare cases such as nuclear chain reactions. In contrast, if $\rho<1$, the system is subcritical, which means that endogenous activity alone is insufficient to sustain itself indefinitely and new exogenous inputs are required to keep the system active. The total expected number of events triggered by a single exogenous event is given by $1/(1-\rho)$ [65]. A system at criticality (i.e., with a spectral radius at or very close to $1$) represents a special regime where a single exogenous event can, in principle, generate an indefinitely long cascade of endogenous events. For example, it is debated whether brain activity [69, 70] and financial transactions operate near criticality [59, 60]. Similarly, research on social contagion and information cascades suggests that systems near criticality optimize responsiveness while avoiding runaway diffusion, enabling the rapid spread of ideas and emotions under the right conditions [71].
Across our 10 bootstrapped estimates, we estimate the spectral radius to be $0.80\pm 0.008$, which places the system well below criticality. This results in an expected cascade size of $5.0\pm 0.2$, meaning that each exogenous emotion, on average, generates a cascade of five total events before activity dies out. Thus, while the system exhibits substantial self-excitation, it remains far from the near-critical regime. A similar pattern holds at the level of individual emotions, where the rate of endogeneity for each emotion type $f$ is given by $\sum_{e}\alpha^{e,f}$.
Zooming in on the entries of the $\alpha^{e,f}$ matrix reveals that each of the 6 emotions displays distinct temporal dynamics (Figure 2, left). One can easily notice that joy has significantly higher levels of self-excitation rate than other emotions, as shown by the comparison between the different diagonal entries $\alpha^{e,e}$ representing self-excitation. Joy is up to 3 times more contagious than even high-arousal negative emotions of anger and disgust. The fact that surprise is the least contagious of all emotions serves as a useful consistency check. Similarly, fear shows relatively low levels of excitation.
Generally, self-exciting emotion contagion effects dominate over emotion interaction effects that are quantified by the off-diagonal values $\alpha^{e,f}$ representing the influence of emotion $f$ on emotion $e$. A notable exception is the moral emotions of anger and disgust, which are mutually reinforcing in almost equal magnitudes. It should also be noted that all emotions, including disgust and sadness, trigger joy. We will come back to these observations in the Discussion section below.
Recall that the decay times $\gamma^{e}$ characterize the direct time scale over which past emotions trigger future ones of type $e$. Figure 3 shows that users have a significantly longer direct memory of negative emotions, in particular high-arousal emotions of anger and disgust. Coupled with the fact that emotions are triggered primarily through self-excitation, this suggests that negative emotions impose a more lasting direct influence on future interactions.
Discussion
Endogenous Influences Dominate Exogenous Influences
We find that live videos, while central to content delivery, have limited direct influence on users’ emotional expressions. Among 92,412 live chat messages on YouTube, emotional dynamics are primarily shaped by user interactions rather than passive video consumption, with peer-driven effects outweighing video influence by a factor of four. This supports the view that emotions are inherently communicative, emerging through social interaction rather than passive reception [72].
The dominance of peer interaction in shaping online emotional expression is consistent with Social Presence Theory [73], which posits that communication channels vary in perceived social presence, influencing interaction patterns. This theory, widely applied in research on online learning [74], suggests that active engagement enhances social presence. In live-stream settings, an interactive audience heightens perceived presence in chat discussions, fostering greater emotional expression than passive video viewing. Specifically, the unique interactive features of live chats encourage social interactions and community forming which are identified as primary drivers for live-stream engagement [75, 76].
Beyond general social influence, emotional mimicry and social appraisal processes further explain the primacy of peer interactions. Emotional mimicry depends on contextual interpretation rather than direct replication of expressions [77]. On YouTube Live, video content provides the contextual backdrop, but emotions propagate mainly through user interactions. Similarly, emotional expressions function as social signals, helping individuals navigate ambiguity by inferring meaning from others’ emotions [78]. In this setting, chat-based emotional expressions guide audience engagement with video content. This aligns with Affect Theory of Social Exchange [79], which holds that emotions emerging from social interactions reinforce group cohesion and collective experiences.
Together, these findings suggest that YouTube Live fosters a socially constructed emotional environment, where peer interactions - not the video itself - drive emotional expression. Consequently, as language models grow more sophisticated, the threat of bots mimicking human users to endogenously manipulate collective emotions becomes an increasing concern [80].
Positive Emotions Are More Contagious than Negative Emotions
Our statistical model allows us to disentangle emotional self- and cross-excitation in live chats. We find that self-excitation clearly dominates: each emotion is most likely to trigger a recurrence of itself rather than give rise to a different emotion. This may help explain why much of the existing research on emotional contagion in online settings has focused on emotions in isolation. Nonetheless, non-negligible cross-excitation effects, discussed below, indicate that interactions between emotions also play a role.
Within self-excitation, we observe a clear hierarchy. Joy exhibits the strongest contagion, followed by moderate levels for anger, disgust, and sadness, and minimal levels for surprise and fear. Specifically, joy’s self-excitation intensity (0.77) is more than three times that of anger and disgust (0.25) and nearly four times that of sadness (0.21). This pattern is consistent with prior findings on users’ tendency to express and amplify positive emotions on social platforms [27, 81], and may be rooted in the evolutionary role of positive emotions in fostering social bonding and cooperation [82]. Anger, disgust, and sadness also show meaningful levels of self-excitation, supporting their function as social signals that coordinate group responses to video content [83, 84].
By contrast, surprise and fear show little evidence of contagion. This is consistent with the fleeting nature of surprise, which often transitions into other emotions following appraisal and sense-making, limiting its transmissibility [85]. Surprise is also shown to be less contagious in face-to-face interactions [86]. Fear, likewise, may show limited contagion due to its strong dependence on situational context [87].
Taken together, this hierarchy of emotional contagion - led by joy - highlights how both the psychological profile and social function of emotions shape their transmissibility in online environments.
Emotions Cross-Excite Each Other
While emotional contagion has been widely studied, less attention has been given to how different emotions trigger one another in large-scale social interactions. Much of the existing literature has focused on how individuals transition between emotional states [35], with some evidence that emotions of similar valence can mutually reinforce one another [36]. In social contexts, one person’s emotional expression may elicit not only mimicry but also different emotional responses in others, depending on the interpretation and context [37]. The notion of emotion cycles has been proposed to describe how emotions within groups recursively shape each other over time [37], and people are known to use multidimensional mental models to predict emotional transitions in others [38].
We contribute to this line of research by analyzing the dynamic interplay of six basic emotions in online social interactions. Our results reveal two prominent modes of cross-excitation.
First, anger and disgust strongly reinforce one another. Their reciprocal triggering, along with high self-excitation, aligns with the role of these moral emotions in driving moral outrage. Previous work has shown that co-occurrence of anger and disgust is predictive of moral outrage [88]; our findings suggest this co-occurrence may be further amplified by mutual excitation. This mechanism could help explain the frequent escalation of outrage in online discourse and its potential contribution to polarization and social fragmentation [89, 90]. We note, however, that part of this effect may be due to the frequent co-labeling of anger and disgust in the same message (see SI Appendix).
Second, we observe a counterintuitive pattern in which negative emotions such as disgust and sadness trigger joy. This may reflect Schadenfreude, trolling and antisocial behaviors, where users express positive affect in response to others’ negative emotions. Online disinhibition and platform affordances may encourage such dynamics, reinforcing gratification from provocation rather than empathy [45, 46, 47]. These findings underscore the need for platform designs that actively encourage prosocial behavior on social media [91].
Positive Emotions are More Contagious, Negative Emotions Last Longer
Research on emotional contagion in online networks has yielded mixed findings on positivity and negativity biases. While self-representation and social validation mechanisms promote positivity [25, 26, 27], negativity often spreads more easily due to cognitive biases that prioritize negative information [28, 29, 30]. Some studies suggest that positive emotions are more contagious [4, 31, 32], whereas others find that negative content tends to reach broader audiences [33, 34], highlighting the context-dependent nature of emotional virality [24].
Many of these studies, however, do not account for differences in the temporal dynamics and base frequencies of emotional expression. By contrast, our Hawkes process framework allows us to disentangle the temporal dependencies governing emotional contagion. Our findings indicate that positivity and negativity biases manifest along distinct dimensions. Positive emotions are more contagious: a single positive message triggers, on average, 0.77 additional positive messages. By comparison, negative emotions such as anger and disgust trigger only 0.26 further negative messages on average.
However, the time scale of these cascades differs markedly. Consistent with Negativity Bias Theory [92], negative emotions unfold over a time window nearly twice as long as that of positive emotions, suggesting that although less contagious in volume, negative affect persists longer in online conversations. These findings resonate with previous observations that positive emotions rise quickly and fade faster, while negative emotions accumulate gradually before dissipating [13, 14, 15].
Limitations
A key limitation of our analysis is that we approximate the emotional content of the video solely through its subtitles. While this text-based approach captures the semantic cues of emotional expression, it omits prosodic and visual signals such as tone of voice, facial expressions, and body language. Previous studies have shown that discrete emotion measures derived from text correlate with self-reported emotions, but may diverge from facial and vocal indicators [48]. Future work integrating multimodal data—including audio and visual cues—could improve emotion inference and offer a more complete picture of how emotional dynamics unfold across communication channels.
Another limitation is that we do not distinguish between human users and automated accounts (bots) in YouTube live chats. This distinction would enable a more refined application of our framework, allowing the analysis of emotion dynamics separately for human and automated agents. Given the increasing prevalence of bots and the rapid development of large language models capable of simulating emotional expression, future studies should examine how emotional exchanges evolve in mixed human-AI environments [93, 94, 95].
Finally, our modeling framework is based on a discrete event Hawkes process that treats emotions as categorical states. While this setup easily generalizes to more than six discrete emotions, it does not directly support alternative representations that view emotions as points in a continuous valence–arousal space [96]. To capture such continuous affective states, the framework could be extended to a marked Hawkes process, where each event carries additional information or marks - such as a real-valued vector indicating valence and arousal. This extension would allow the model to capture not only the timing of events but also their qualitative characteristics, making it well suited to reflect the graded nature of emotional responses in continuous spaces.
Conclusions
We analyzed 92,412 YouTube Live chat messages from 397 videos to study emotional contagion in social networks and livestream platforms. Our findings show that user interactions are four times more emotionally influential than passive video content. Joy is three times more contagious than negative emotions, but negative emotions persist nearly twice as long, potentially creating lasting negative moods. We also observed cross-excitation effects, where negative posts can trigger positive emotions in trolls, revealing antisocial dynamics that harm community cohesion. These insights have practical implications: promoting joyful content and enabling active communication (e.g., live chats or bots) can foster emotional engagement and positivity. However, the lingering impact of negative emotions highlights the need for timely moderation. Platforms must balance emotional amplification with community well-being. Designing for positive contagion while managing toxic interactions is key to sustaining healthy online environments. Future research could explore cultural influences and long-term impacts of emotional contagion strategies.
@@ Appendix A Summary Statistics of YouTube Live Chat Data
We show some additional summary statistics of our dataset. Once we filtered the data by removing inactive videos, we compiled 92,412 live chat messages from 397 videos, amassing a total of 780 hours of material and encompassing 27 distinct topics. The distribution of topics is shown in Figure A.1. The distribution of emotional events is shown in Figure A.2. We also show the distribution of the median time between each event arrival in Figure A.3. Since we label emotions non-exclusively, we show the co-occurrence patterns across 6 basic emotions in Figure A.4. In general, negative emotions have a higher tendency to co-occur with one another. In particular, anger and disgust have a relatively high co-occurrence.
@@ Appendix B Robustness Analysis on Video Influence
To examine the impact of video content on live chat discussions, we have parameterized and interpolated the prevalence of emotions in the video via sums of log-normal shapes (see Methods). Here, we test the robustness of our results with respect to changes in the log-normal parameters. We have defined the shape of the log-normal function by calibrating its mode (maximum value), and the median. We assumed that the intensity of each transcript peaks $2$ seconds after appearance such that the maximum of the log-normal function lies at $2$ seconds. We further assumed that 50% of the emotion intensity for each transcript is manifested within $10$ seconds of transcript appearance such that the median is equal to $10$ seconds. Here we repeat our analysis while changing the log-normal function maximum and median by factors of 0.5 and 2 to test the robustness of our results. This effectively changes assumptions about the audience’s reaction time and memory of prior content. We observe in Figures 1(a), 1(b), 2(a), and 2(b) that the results are qualitatively consistent with the results in the main text. This suggests that our results aren’t particularly sensitive to the parametrization of the video content.
@@ Appendix C Dynamics of 4 Emotions
Our main analysis models the commonly studied 6 basic emotions [50]. Here we test the robustness of our results by modeling $|\mathcal{E}|=4$ emotions: joy, anger, disgust, and sadness, removing fear and surprise because they have small excitatory patterns and low overall intensity. We obtain a sample of 288,581 live chat messages across 1,405 videos. Reducing the dimensionality of the multivariate Hawkes process reduces the number of parameters to be estimated from 84 to 40. We obtain our results for 4 emotions in Figures C.1 and C.2. We find that the dynamics of 4 emotions are qualitatively similar to our main results on 6 emotions (Figures 2 and 3 in the main text).
@@ Appendix E Dynamics of Live Chats with No Emotional Content
Table E.1: The mean and standard deviation of estimated parameters for a univariate Hawkes Process of live chat messages with no emotion labels across 10 bootstrapped samples.
@@
Our main study shows that online emotions are characterized by excitatory patterns. Despite the constant feed of video content, the emotion dynamics evolve mostly spontaneously among users. To complement our analysis, we study the behavior of live chats without emotional content. We train a transformer to assign emotion labels for live chats. We consider live chat messages that were not assigned any emotion labels to contain no emotional content. This gives us 66,235 live chat messages with no emotional content across 232 videos. We use a univariate Hawkes process to describe the intensity of live chat messages without emotion labels. The intensity $\lambda(t)$ at time $t$ during a given video is specified as
@@
The value of $S(t)$ is given by $S(t)=\sum_{\tau_{j}<t}s_{\tau_{j}}(t)$ where $\left\{\tau_{j}\right\}$ enumerates all times at which a subtitle first appears in the video. The values of $s_{\tau_{j}}(t)$ are calculated according to Equation (3) from the main text with $\mu=2.3$ and $\sigma=1.3$. The term $s_{\tau_{j}}(t)$ represents the time-varying intensity arising from a video subscript appearing in the video at time $\tau_{j}$. Similarly, we assume that $\phi$ is exponentially decaying where $\phi(t)=\alpha\frac{1}{\gamma}e^{-\frac{1}{\gamma}t}$ with decay time $\gamma$. Table E.1 shows the estimated parameters of Equation (1). We observe that live chat messages with no emotional content are nonetheless highly self-exciting but with a shorter direct memory than emotional live chats. Unlike our main result (Figure 3 in the main text), live chat messages with no emotional content are driven more equally by the video and user spontaneity, as shown in Figure E.1. This shows that while live discussions evolve largely endogenously, live chats with no emotional content are much more affected by videos. In contrast, emotions on online platforms are largely attributed to user expressions and interactions.